Assets

Liabilities

Reserves                    $ 1,500

Loans                       $ 3,500

Initial Deposit           $ 5,000

Assume required reserve ratio is 10 percent and that 100% of all the loans are deposited as new deposits.


                 

               ii.           Calculate the deposit and credit multiplier.

              

                


1
Expert's answer
2021-06-21T12:05:15-0400

Given,

Initial deposits = $5000

Required reserve ratio = 10 percent

 

Required reserve ratio is the part of the total deposit that the bank has to always keep with themselves. They cannot lend that money.

Therefore required reserve would be = 10% of 5000

                required reserve"= \frac{10}{100} \times 5000"

                required reserve = 500

The bank could lend = $4500

Total deposits now would be"= 5000 + 4500"

                        "= \$9500"


Credit multiplier is calculated using the following formula:

Credit multiplier = Change in deposits"\div" Change in reserve

Credit multiplier"= \frac{4500}{500}"

Credit multiplier = 9